OpenStudy (anonymous):

Someone please help me out with this! Solving Linear Systems by Using Graphs and Tables - Algebra 2 Use a graph and a table to solve the system shown below. Be sure to check your answer and to enter it like this example: (5, -3). {3y + 6x = 3 {x - y = -7 Could someone show me how to do this, please?

OpenStudy (mathmale):

You are to create two tables of x and y values, one for each of the given linear equations. For the first equation, 3y+6x=3: Solve for y. Result: 3y=3-6x, or y=1-2x. Make a table, as show. Choose x values at random and calculate y for each one.